. A wooden ramp has a coefficient of static friction of 0.408 and an angle of inclination of 36.2°. A 4.74 kg book is placed on the ramp. Does the book move? Show why or why not.


1
Expert's answer
2021-01-28T19:38:20-0500

Let's apply the Newtons Second Law of Motion in projections on axis "x" and "y" :


"F_{fr}-mgsin\theta=0,""N-mgcos\theta=0."


If the horizontal component of the book's weight will be greater than the force of friction, the book will move. Let's check that:


"F_{fr}=\mu mgcos\theta,""F_{fr}=0.408\cdot4.74\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot cos36.2^{\circ}=15.29\ N,""W_{||}=mgsin\theta=4.74\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot sin36.2^{\circ}=27.43\ N."

As we can see from the calculations, "W_{||}>F_{fr}". Therefore, the book will move.